OpenStudy (anonymous):

Before 1976, California license-plate numbers consisted of three letters followed by three numbers (for example, EGT 893). In the summer of 1976, however, the state ran out of numbers of this type and started issuing seven-digit plates (such as GIR 788 1). How many possible license plate numbers can be made from 3 letters followed by 4 numerical digits? (Base your answer on 26 letters in the alphabet and 10 numerical digits.)

OpenStudy (anonymous):

A. 17,576,000 B. 17,576 C. 1,000,000 D. 175,760,000 E. 10,000,000

OpenStudy (anonymous):

A? Okay thanks!

ganeshie8 (ganeshie8):

26 alphabets and 10 digits to pick from. \(\huge \textbf{_ _ _} \ \textbf{ _ _ _ _} \) you can fill first 3 spaces with alphabets, last 4 with digits 26*26*26 * 10*10*10*10 = 26^3 * 10^4

OpenStudy (anonymous):

26^3* ( 10^4) = option D

OpenStudy (anonymous):

So it's D? haha

ganeshie8 (ganeshie8):

yep! only D has four 0 s in the end :D

OpenStudy (anonymous):

Thanks !!!!